# Fairness-Aware Domain Generalization under Covariate and Dependence Shifts

### v1

@@ Appendix F Additional Results
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (B, 0.91) (E, 0.87) (I, 0.58)
RandAug 0.64$\pm$0.26 / 0.64$\pm$0.15 / 93.47$\pm$1.56 0.41$\pm$0.34 / 0.68$\pm$0.09 / 95.62$\pm$1.96 0.44$\pm$0.21 / 0.63$\pm$0.05 / 92.99$\pm$1.00
ERM 0.67$\pm$0.17 / 0.58$\pm$0.02 / 91.89$\pm$1.10 0.43$\pm$0.21 / 0.64$\pm$0.02 / 95.69$\pm$2.19 0.50$\pm$0.19 / 0.59$\pm$0.03 / 93.28$\pm$1.61
IRM 0.63$\pm$0.12 / 0.58$\pm$0.01 / 93.39$\pm$1.03 0.32$\pm$0.23 / 0.63$\pm$0.03 / 95.12$\pm$0.49 0.45$\pm$0.06 / 0.59$\pm$0.02 / 92.01$\pm$1.13
GDRO 0.71$\pm$0.16 / 0.57$\pm$0.02 / 89.81$\pm$1.10 0.46$\pm$0.16 / 0.61$\pm$0.02 / 95.26$\pm$1.53 0.50$\pm$0.14 / 0.59$\pm$0.01 / 93.27$\pm$1.27
Mixup 0.58$\pm$0.19 / 0.59$\pm$0.02 / 92.46$\pm$0.69 0.40$\pm$0.04 / 0.61$\pm$0.02 / 93.31$\pm$1.42 0.42$\pm$0.09 / 0.59$\pm$0.02 / 93.42$\pm$2.43
MLDG 0.63$\pm$0.25 / 0.58$\pm$0.02 / 92.71$\pm$2.36 0.41$\pm$0.15 / 0.62$\pm$0.03 / 95.59$\pm$0.87 0.51$\pm$0.15 / 0.60$\pm$0.02 / 93.35$\pm$1.87
CORAL 0.69$\pm$0.19 / 0.58$\pm$0.01 / 92.09$\pm$2.03 0.34$\pm$0.24 / 0.64$\pm$0.01 / 95.91$\pm$1.44 0.53$\pm$0.05 / 0.59$\pm$0.02 / 93.35$\pm$0.26
MMD 0.69$\pm$0.25 / 0.56$\pm$0.01 / 93.87$\pm$0.14 0.45$\pm$0.22 / 0.57$\pm$0.02 / 94.68$\pm$0.20 0.27$\pm$0.18 / 0.57$\pm$0.03 / 89.88$\pm$0.22
DANN 0.46$\pm$0.07 / 0.61$\pm$0.02 / 91.80$\pm$0.64 0.53$\pm$0.18 / 0.85$\pm$0.03 / 91.54$\pm$2.24 0.38$\pm$0.18 / 0.63$\pm$0.01 / 90.09$\pm$0.60
CDANN 0.62$\pm$0.24 / 0.59$\pm$0.03 / 91.22$\pm$0.33 0.43$\pm$0.10 / 0.66$\pm$0.02 / 94.75$\pm$2.23 0.43$\pm$0.18 / 0.61$\pm$0.01 / 92.41$\pm$1.68
DDG 0.60$\pm$0.20 / 0.59$\pm$0.02 / 91.76$\pm$1.03 0.36$\pm$0.15 / 0.63$\pm$0.02 / 95.52$\pm$2.35 0.49$\pm$0.17 / 0.59$\pm$0.01 / 92.35$\pm$2.04
MBDG 0.60$\pm$0.15 / 0.58$\pm$0.01 / 91.29$\pm$1.41 0.46$\pm$0.19 / 0.63$\pm$0.01 / 95.01$\pm$1.39 0.52$\pm$0.14 / 0.58$\pm$0.02 / 92.77$\pm$2.07
DDG-FC 0.61$\pm$0.06 / 0.58$\pm$0.03 / 92.27$\pm$1.65 0.39$\pm$0.18 / 0.64$\pm$0.03 / 95.51$\pm$2.36 0.45$\pm$0.17 / 0.58$\pm$0.03 / 93.38$\pm$0.52
MBDG-FC 0.70$\pm$0.15 / 0.56$\pm$0.03 / 92.12$\pm$0.43 0.35$\pm$0.07 / 0.60$\pm$0.01 / 95.54$\pm$1.80 0.56$\pm$0.07 / 0.57$\pm$0.01 / 92.41$\pm$1.61
EIIL 0.88$\pm$0.07 / 0.59$\pm$0.05 / 84.75$\pm$2.16 0.69$\pm$0.12 / 0.71$\pm$0.01 / 92.86$\pm$1.70 0.47$\pm$0.08 / 0.57$\pm$0.01 / 86.93$\pm$0.89
FarconVAE 0.93$\pm$0.03 / 0.54$\pm$0.01 / 89.61$\pm$0.64 0.72$\pm$0.17 / 0.63$\pm$0.01 / 91.50$\pm$1.89 0.42$\pm$0.24 / 0.58$\pm$0.03 / 87.42$\pm$2.14
FATDM 0.93$\pm$0.03 / 0.57$\pm$0.02 / 92.20$\pm$0.36 0.80$\pm$0.02 / 0.65$\pm$0.02 / 92.89$\pm$1.00 0.52$\pm$0.10 / 0.60$\pm$0.01 / 92.22$\pm$1.60
FDDG 0.94$\pm$0.05 / 0.55$\pm$0.02 / 93.91$\pm$0.33 0.87$\pm$0.05 / 0.60$\pm$0.01 / 95.91$\pm$1.06 0.48$\pm$0.06 / 0.57$\pm$0.02 / 92.55$\pm$1.45
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (M, 0.87) (S, 0.39) (W, 0.49)
RandAug 0.36$\pm$0.12 / 0.65$\pm$0.05 / 92.79$\pm$1.22 0.35$\pm$0.20 / 0.69$\pm$0.06 / 91.89$\pm$1.02 0.34$\pm$0.09 / 0.64$\pm$0.02 / 92.07$\pm$0.55
ERM 0.34$\pm$0.08 / 0.62$\pm$0.01 / 92.51$\pm$1.45 0.68$\pm$0.14 / 0.59$\pm$0.03 / 93.48$\pm$0.94 0.39$\pm$0.09 / 0.61$\pm$0.01 / 92.82$\pm$0.38
IRM 0.34$\pm$0.11 / 0.65$\pm$0.02 / 92.47$\pm$2.42 0.55$\pm$0.23 / 0.59$\pm$0.01 / 91.81$\pm$0.66 0.32$\pm$0.19 / 0.66$\pm$0.01 / 90.54$\pm$1.56
GDRO 0.45$\pm$0.14 / 0.63$\pm$0.02 / 91.75$\pm$1.11 0.72$\pm$0.14 / 0.59$\pm$0.01 / 93.65$\pm$0.67 0.48$\pm$0.09 / 0.60$\pm$0.01 / 92.50$\pm$0.38
Mixup 0.31$\pm$0.11 / 0.62$\pm$0.02 / 93.52$\pm$0.79 0.91$\pm$0.04 / 0.58$\pm$0.02 / 93.20$\pm$0.33 0.43$\pm$0.19 / 0.61$\pm$0.01 / 92.98$\pm$0.03
MLDG 0.35$\pm$0.20 / 0.62$\pm$0.01 / 92.45$\pm$0.07 0.71$\pm$0.22 / 0.57$\pm$0.01 / 93.85$\pm$0.40 0.47$\pm$0.20 / 0.59$\pm$0.01 / 92.82$\pm$1.65
CORAL 0.43$\pm$0.08 / 0.63$\pm$0.01 / 92.23$\pm$0.06 0.74$\pm$0.10 / 0.58$\pm$0.01 / 93.77$\pm$1.99 0.50$\pm$0.14 / 0.60$\pm$0.02 / 92.47$\pm$2.04
MMD 0.48$\pm$0.25 / 0.62$\pm$0.02 / 91.07$\pm$2.00 0.66$\pm$0.18 / 0.59$\pm$0.03 / 92.58$\pm$1.63 0.39$\pm$0.20 / 0.68$\pm$0.02 / 91.75$\pm$1.37
DANN 0.65$\pm$0.14 / 0.88$\pm$0.01 / 91.46$\pm$0.50 0.80$\pm$0.14 / 0.57$\pm$0.02 / 88.20$\pm$1.65 0.11$\pm$0.09 / 0.66$\pm$0.01 / 86.80$\pm$1.18
CDANN 0.27$\pm$0.12 / 0.67$\pm$0.01 / 91.07$\pm$0.97 0.52$\pm$0.12 / 0.82$\pm$0.02 / 88.32$\pm$0.37 0.35$\pm$0.17 / 0.67$\pm$0.02 / 90.19$\pm$0.60
DDG 0.37$\pm$0.14 / 0.64$\pm$0.01 / 91.36$\pm$0.65 0.63$\pm$0.22 / 0.58$\pm$0.01 / 93.40$\pm$0.37 0.51$\pm$0.07 / 0.60$\pm$0.01 / 91.34$\pm$0.80
MBDG 0.38$\pm$0.14 / 0.64$\pm$0.02 / 92.23$\pm$1.15 0.67$\pm$0.06 / 0.56$\pm$0.03 / 93.12$\pm$0.70 0.30$\pm$0.04 / 0.62$\pm$0.01 / 91.05$\pm$0.53
DDG-FC 0.42$\pm$0.09 / 0.95$\pm$0.03 / 92.70$\pm$1.49 0.76$\pm$0.21 / 0.59$\pm$0.02 / 93.85$\pm$1.79 0.48$\pm$0.15 / 0.62$\pm$0.02 / 92.45$\pm$1.55
MBDG-FC 0.49$\pm$0.19 / 0.63$\pm$0.03 / 90.67$\pm$0.42 0.74$\pm$0.23 / 0.57$\pm$0.01 / 93.24$\pm$0.32 0.32$\pm$0.07 / 0.60$\pm$0.03 / 91.50$\pm$0.57
EIIL 0.52$\pm$0.09 / 0.63$\pm$0.03 / 84.96$\pm$1.37 0.98$\pm$0.01 / 0.55$\pm$0.02 / 89.99$\pm$2.27 0.46$\pm$0.05 / 0.65$\pm$0.03 / 86.53$\pm$1.02
FarconVAE 0.54$\pm$0.22 / 0.58$\pm$0.02 / 85.62$\pm$1.49 0.92$\pm$0.06 / 0.56$\pm$0.10 / 90.00$\pm$0.05 0.51$\pm$0.07 / 0.60$\pm$0.01 / 86.40$\pm$0.42
FATDM 0.55$\pm$0.12 / 0.65$\pm$0.01 / 92.23$\pm$1.56 0.92$\pm$0.10 / 0.57$\pm$0.02 / 92.36$\pm$0.99 0.46$\pm$0.05 / 0.63$\pm$0.01 / 92.56$\pm$0.31
FDDG 0.54$\pm$0.08 / 0.62$\pm$0.02 / 92.61$\pm$1.84 0.98$\pm$0.01 / 0.55$\pm$0.01 / 92.26$\pm$2.48 0.52$\pm$0.17 / 0.58$\pm$0.03 / 93.02$\pm$0.50
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (L, 0.48) Avg
RandAug 0.39$\pm$0.10 / 0.70$\pm$0.02 / 91.77$\pm$0.61 0.42 / 0.66 / 92.94
ERM 0.57$\pm$0.15 / 0.62$\pm$0.01 / 91.96$\pm$0.51 0.51 / 0.61 / 93.08
IRM 0.41$\pm$.021 / 0.63$\pm$0.05 / 92.06$\pm$1.89 0.43 / 0.62 / 92.48
GDRO 0.54$\pm$0.15 / 0.62$\pm$0.01 / 91.59$\pm$0.51 0.55 / 0.60 / 92.55
Mixup 0.55$\pm$0.22 / 0.61$\pm$0.02 / 93.43$\pm$2.02 0.51 / 0.60 / 93.19
MLDG 0.53$\pm$0.18 / 0.62$\pm$0.03 / 92.99$\pm$0.86 0.51 / 0.60 / 93.39
CORAL 0.56$\pm$0.23 / 0.59$\pm$0.03 / 92.62$\pm$1.11 0.54 / 0.60 / 93.21
MMD 0.55$\pm$0.16 / 0.61$\pm$0.02 / 92.53$\pm$1.41 0.50 / 0.60 / 92.34
DANN 0.39$\pm$0.21 / 0.67$\pm$0.01 / 90.82$\pm$2.44 0.47 / 0.70 / 90.10
CDANN 0.42$\pm$0.23 / 0.61$\pm$0.03 / 92.42$\pm$2.19 0.43 / 0.66 / 91.48
DDG 0.44$\pm$0.17 / 0.62$\pm$0.02 / 93.46$\pm$0.32 0.49 / 0.61 / 92.74
MBDG 0.56$\pm$0.09 / 0.61$\pm$0.01 / 93.49$\pm$0.97 0.50 / 0.60 / 92.71
DDG-FC 0.50$\pm$0.25 / 0.62$\pm$0.03 / 92.42$\pm$0.30 0.52 / 0.61 / 93.23
MBDG-FC 0.57$\pm$0.23 / 0.62$\pm$0.02 / 91.89$\pm$0.81 0.53 / 0.60 / 92.48
EIIL 0.49$\pm$0.07 / 0.59$\pm$0.01 / 88.39$\pm$1.25 0.64 / 0.61 / 87.78
FarconVAE 0.58$\pm$0.05 / 0.60$\pm$0.05 / 88.70$\pm$0.71 0.66 / 0.58 / 88.46
FATDM 0.51$\pm$0.16 / 0.63$\pm$0.02 / 93.33$\pm$0.20 0.67 / 0.61 / 92.54
FDDG 0.58$\pm$0.15 / 0.59$\pm$0.01 / 93.73$\pm$0.26 0.70 / 0.58 / 93.42
@@
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$

### v2

@@ Appendix E Additional Results
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
Methods (B, 0.91) (E, 0.87) (I, 0.58) (W, 0.49) (L, 0.48)
ColorJitter 0.64$\pm$0.26 / 0.64$\pm$0.15 / 93.47$\pm$1.56 0.41$\pm$0.34 / 0.68$\pm$0.09 / 95.62$\pm$1.96 0.44$\pm$0.21 / 0.63$\pm$0.05 / 92.99$\pm$1.00 0.34$\pm$0.09 / 0.64$\pm$0.02 / 92.07$\pm$0.55 0.39$\pm$0.10 / 0.70$\pm$0.02 / 91.77$\pm$0.61
ERM 0.67$\pm$0.17 / 0.58$\pm$0.02 / 91.89$\pm$1.10 0.43$\pm$0.21 / 0.64$\pm$0.02 / 95.69$\pm$2.19 0.50$\pm$0.19 / 0.59$\pm$0.03 / 93.28$\pm$1.61 0.39$\pm$0.09 / 0.61$\pm$0.01 / 92.82$\pm$0.38 0.57$\pm$0.15 / 0.62$\pm$0.01 / 91.96$\pm$0.51
IRM 0.63$\pm$0.12 / 0.58$\pm$0.01 / 93.39$\pm$1.03 0.32$\pm$0.23 / 0.63$\pm$0.03 / 95.12$\pm$0.49 0.45$\pm$0.06 / 0.59$\pm$0.02 / 92.01$\pm$1.13 0.32$\pm$0.19 / 0.66$\pm$0.01 / 90.54$\pm$1.56 0.41$\pm$.021 / 0.63$\pm$0.05 / 92.06$\pm$1.89
GDRO 0.71$\pm$0.16 / 0.57$\pm$0.02 / 89.81$\pm$1.10 0.46$\pm$0.16 / 0.61$\pm$0.02 / 95.26$\pm$1.53 0.50$\pm$0.14 / 0.59$\pm$0.01 / 93.27$\pm$1.27 0.48$\pm$0.09 / 0.60$\pm$0.01 / 92.50$\pm$0.38 0.54$\pm$0.15 / 0.62$\pm$0.01 / 91.59$\pm$0.51
Mixup 0.58$\pm$0.19 / 0.59$\pm$0.02 / 92.46$\pm$0.69 0.40$\pm$0.04 / 0.61$\pm$0.02 / 93.31$\pm$1.42 0.42$\pm$0.09 / 0.59$\pm$0.02 / 93.42$\pm$2.43 0.43$\pm$0.19 / 0.61$\pm$0.01 / 92.98$\pm$0.03 0.55$\pm$0.22 / 0.61$\pm$0.02 / 93.43$\pm$2.02
MLDG 0.63$\pm$0.25 / 0.58$\pm$0.02 / 92.71$\pm$2.36 0.41$\pm$0.15 / 0.62$\pm$0.03 / 95.59$\pm$0.87 0.51$\pm$0.15 / 0.60$\pm$0.02 / 93.35$\pm$1.87 0.47$\pm$0.20 / 0.59$\pm$0.01 / 92.82$\pm$1.65 0.53$\pm$0.18 / 0.62$\pm$0.03 / 92.99$\pm$0.86
CORAL 0.69$\pm$0.19 / 0.58$\pm$0.01 / 92.09$\pm$2.03 0.34$\pm$0.24 / 0.64$\pm$0.01 / 95.91$\pm$1.44 0.53$\pm$0.05 / 0.59$\pm$0.02 / 93.35$\pm$0.26 0.50$\pm$0.14 / 0.60$\pm$0.02 / 92.47$\pm$2.04 0.56$\pm$0.23 / 0.59$\pm$0.03 / 92.62$\pm$1.11
MMD 0.69$\pm$0.25 / 0.56$\pm$0.01 / 93.87$\pm$0.14 0.45$\pm$0.22 / 0.57$\pm$0.02 / 94.68$\pm$0.20 0.27$\pm$0.18 / 0.57$\pm$0.03 / 89.88$\pm$0.22 0.39$\pm$0.20 / 0.68$\pm$0.02 / 91.75$\pm$1.37 0.55$\pm$0.16 / 0.61$\pm$0.02 / 92.53$\pm$1.41
DANN 0.46$\pm$0.07 / 0.61$\pm$0.02 / 91.80$\pm$0.64 0.53$\pm$0.18 / 0.85$\pm$0.03 / 91.54$\pm$2.24 0.38$\pm$0.18 / 0.63$\pm$0.01 / 90.09$\pm$0.60 0.11$\pm$0.09 / 0.66$\pm$0.01 / 86.80$\pm$1.18 0.39$\pm$0.21 / 0.67$\pm$0.01 / 90.82$\pm$2.44
CDANN 0.62$\pm$0.24 / 0.59$\pm$0.03 / 91.22$\pm$0.33 0.43$\pm$0.10 / 0.66$\pm$0.02 / 94.75$\pm$2.23 0.43$\pm$0.18 / 0.61$\pm$0.01 / 92.41$\pm$1.68 0.35$\pm$0.17 / 0.67$\pm$0.02 / 90.19$\pm$0.60 0.42$\pm$0.23 / 0.61$\pm$0.03 / 92.42$\pm$2.19
DDG 0.60$\pm$0.20 / 0.59$\pm$0.02 / 91.76$\pm$1.03 0.36$\pm$0.15 / 0.63$\pm$0.02 / 95.52$\pm$2.35 0.49$\pm$0.17 / 0.59$\pm$0.01 / 92.35$\pm$2.04 0.51$\pm$0.07 / 0.60$\pm$0.01 / 91.34$\pm$0.80 0.44$\pm$0.17 / 0.62$\pm$0.02 / 93.46$\pm$0.32
MBDG 0.60$\pm$0.15 / 0.58$\pm$0.01 / 91.29$\pm$1.41 0.46$\pm$0.19 / 0.63$\pm$0.01 / 95.01$\pm$1.39 0.52$\pm$0.14 / 0.58$\pm$0.02 / 92.77$\pm$2.07 0.30$\pm$0.04 / 0.62$\pm$0.01 / 91.05$\pm$0.53 0.56$\pm$0.09 / 0.61$\pm$0.01 / 93.49$\pm$0.97
DDG-FC 0.61$\pm$0.06 / 0.58$\pm$0.03 / 92.27$\pm$1.65 0.39$\pm$0.18 / 0.64$\pm$0.03 / 95.51$\pm$2.36 0.45$\pm$0.17 / 0.58$\pm$0.03 / 93.38$\pm$0.52 0.48$\pm$0.15 / 0.62$\pm$0.02 / 92.45$\pm$1.55 0.50$\pm$0.25 / 0.62$\pm$0.03 / 92.42$\pm$0.30
MBDG-FC 0.70$\pm$0.15 / 0.56$\pm$0.03 / 92.12$\pm$0.43 0.35$\pm$0.07 / 0.60$\pm$0.01 / 95.54$\pm$1.80 0.56$\pm$0.07 / 0.57$\pm$0.01 / 92.41$\pm$1.61 0.32$\pm$0.07 / 0.60$\pm$0.03 / 91.50$\pm$0.57 0.57$\pm$0.23 / 0.62$\pm$0.02 / 91.89$\pm$0.81
EIIL 0.88$\pm$0.07 / 0.59$\pm$0.05 / 84.75$\pm$2.16 0.69$\pm$0.12 / 0.71$\pm$0.01 / 92.86$\pm$1.70 0.47$\pm$0.08 / 0.57$\pm$0.01 / 86.93$\pm$0.89 0.46$\pm$0.05 / 0.65$\pm$0.03 / 86.53$\pm$1.02 0.49$\pm$0.07 / 0.59$\pm$0.01 / 88.39$\pm$1.25
FarconVAE 0.93$\pm$0.03 / 0.54$\pm$0.01 / 89.61$\pm$0.64 0.72$\pm$0.17 / 0.63$\pm$0.01 / 91.50$\pm$1.89 0.42$\pm$0.24 / 0.58$\pm$0.03 / 87.42$\pm$2.14 0.51$\pm$0.07 / 0.60$\pm$0.01 / 86.40$\pm$0.42 0.58$\pm$0.05 / 0.60$\pm$0.05 / 88.70$\pm$0.71
FCR 0.81$\pm$0.05 / 0.59$\pm$0.02 / 79.66$\pm$0.25 0.60$\pm$0.09 / 0.69$\pm$0.02 / 89.22$\pm$1.30 0.40$\pm$0.06 / 0.62$\pm$0.02 / 79.15$\pm$0.56 0.39$\pm$0.06 / 0.63$\pm$0.02 / 82.33$\pm$0.89 0.38$\pm$0.12 / 0.66$\pm$0.02 / 85.22$\pm$2.33
FTCS 0.75$\pm$0.10 / 0.60$\pm$0.02 / 80.00$\pm$0.20 0.66$\pm$0.18 / 0.65$\pm$0.01 / 88.11$\pm$1.09 0.49$\pm$0.05 / 0.65$\pm$0.01 / 82.15$\pm$0.64 0.40$\pm$0.06 / 0.60$\pm$0.02 / 79.66$\pm$1.05 0.42$\pm$0.23 / 0.65$\pm$0.03 / 79.64$\pm$1.00
FATDM 0.93$\pm$0.03 / 0.57$\pm$0.02 / 92.20$\pm$0.36 0.80$\pm$0.02 / 0.65$\pm$0.02 / 92.89$\pm$1.00 0.52$\pm$0.10 / 0.60$\pm$0.01 / 92.22$\pm$1.60 0.46$\pm$0.05 / 0.63$\pm$0.01 / 92.56$\pm$0.31 0.51$\pm$0.16 / 0.63$\pm$0.02 / 93.33$\pm$0.20
FEDORA 0.94$\pm$0.05 / 0.55$\pm$0.02 / 93.91$\pm$0.33 0.87$\pm$0.05 / 0.60$\pm$0.01 / 95.91$\pm$1.06 0.48$\pm$0.06 / 0.57$\pm$0.02 / 92.55$\pm$1.45 0.52$\pm$0.17 / 0.58$\pm$0.03 / 93.02$\pm$0.50 0.58$\pm$0.15 / 0.59$\pm$0.01 / 93.73$\pm$0.26
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
Methods (M, 0.87) (S, 0.39) Avg
ColorJitter 0.36$\pm$0.12 / 0.65$\pm$0.05 / 92.79$\pm$1.22 0.35$\pm$0.20 / 0.69$\pm$0.06 / 91.89$\pm$1.02 0.42 / 0.66 / 92.94
ERM 0.34$\pm$0.08 / 0.62$\pm$0.01 / 92.51$\pm$1.45 0.68$\pm$0.14 / 0.59$\pm$0.03 / 93.48$\pm$0.94 0.51 / 0.61 / 93.08
IRM 0.34$\pm$0.11 / 0.65$\pm$0.02 / 92.47$\pm$2.42 0.55$\pm$0.23 / 0.59$\pm$0.01 / 91.81$\pm$0.66 0.43 / 0.62 / 92.48
GDRO 0.45$\pm$0.14 / 0.63$\pm$0.02 / 91.75$\pm$1.11 0.72$\pm$0.14 / 0.59$\pm$0.01 / 93.65$\pm$0.67 0.55 / 0.60 / 92.55
Mixup 0.31$\pm$0.11 / 0.62$\pm$0.02 / 93.52$\pm$0.79 0.91$\pm$0.04 / 0.58$\pm$0.02 / 93.20$\pm$0.33 0.51 / 0.60 / 93.19
MLDG 0.35$\pm$0.20 / 0.62$\pm$0.01 / 92.45$\pm$0.07 0.71$\pm$0.22 / 0.57$\pm$0.01 / 93.85$\pm$0.40 0.51 / 0.60 / 93.39
CORAL 0.43$\pm$0.08 / 0.63$\pm$0.01 / 92.23$\pm$0.06 0.74$\pm$0.10 / 0.58$\pm$0.01 / 93.77$\pm$1.99 0.54 / 0.60 / 93.21
MMD 0.48$\pm$0.25 / 0.62$\pm$0.02 / 91.07$\pm$2.00 0.66$\pm$0.18 / 0.59$\pm$0.03 / 92.58$\pm$1.63 0.50 / 0.60 / 92.34
DANN 0.65$\pm$0.14 / 0.88$\pm$0.01 / 91.46$\pm$0.50 0.80$\pm$0.14 / 0.57$\pm$0.02 / 88.20$\pm$1.65 0.47 / 0.70 / 90.10
CDANN 0.27$\pm$0.12 / 0.67$\pm$0.01 / 91.07$\pm$0.97 0.52$\pm$0.12 / 0.82$\pm$0.02 / 88.32$\pm$0.37 0.43 / 0.66 / 91.48
DDG 0.37$\pm$0.14 / 0.64$\pm$0.01 / 91.36$\pm$0.65 0.63$\pm$0.22 / 0.58$\pm$0.01 / 93.40$\pm$0.37 0.49 / 0.61 / 92.74
MBDG 0.38$\pm$0.14 / 0.64$\pm$0.02 / 92.23$\pm$1.15 0.67$\pm$0.06 / 0.56$\pm$0.03 / 93.12$\pm$0.70 0.50 / 0.60 / 92.71
DDG-FC 0.42$\pm$0.09 / 0.95$\pm$0.03 / 92.70$\pm$1.49 0.76$\pm$0.21 / 0.59$\pm$0.02 / 93.85$\pm$1.79 0.52 / 0.61 / 93.23
MBDG-FC 0.49$\pm$0.19 / 0.63$\pm$0.03 / 90.67$\pm$0.42 0.74$\pm$0.23 / 0.57$\pm$0.01 / 93.24$\pm$0.32 0.53 / 0.60 / 92.48
EIIL 0.52$\pm$0.09 / 0.63$\pm$0.03 / 84.96$\pm$1.37 0.98$\pm$0.01 / 0.55$\pm$0.02 / 89.99$\pm$2.27 0.64 / 0.61 / 87.78
FarconVAE 0.54$\pm$0.22 / 0.58$\pm$0.02 / 85.62$\pm$1.49 0.92$\pm$0.06 / 0.56$\pm$0.10 / 90.00$\pm$0.05 0.66 / 0.58 / 88.46
FRC 0.51$\pm$0.08 / 0.66$\pm$0.02 / 82.16$\pm$0.78 0.72$\pm$0.10 / 0.60$\pm$0.01 / 88.01$\pm$1.00 0.54 / 0.63 / 83.68
FTCS 0.49$\pm$0.10 / 0.68$\pm$0.01 / 81.15$\pm$1.25 0.75$\pm$0.21 / 0.62$\pm$0.02 / 75.69$\pm$2.07 0.57 / 0.64 / 80.91
FATDM 0.55$\pm$0.12 / 0.65$\pm$0.01 / 92.23$\pm$1.56 0.92$\pm$0.10 / 0.57$\pm$0.02 / 92.36$\pm$0.99 0.67 / 0.61 / 92.54
FEDORA 0.54$\pm$0.08 / 0.62$\pm$0.02 / 92.61$\pm$1.84 0.98$\pm$0.01 / 0.55$\pm$0.01 / 92.26$\pm$2.48 0.70 / 0.58 / 93.42
@@
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$